OpenStudy (anonymous):

write an equation in standard form with integer coefficients for the line with slope 13/2 going through the point (-3,-4).

Directrix (directrix):

Start with point-slope form of a line y - y1 = m (x - x1)

OpenStudy (anonymous):

can you go step by step?

Directrix (directrix):

y - y1 = (13/2) (x - x1) y - (-4) = (13/2) ( x - (-3))

Directrix (directrix):

y + 4 = (13/2) ( x + 3)

OpenStudy (anonymous):

?

Directrix (directrix):

Multiply through by 2 2(y + 4) = 2 [(13/2) ( x + 3)] 2(y + 4) = (13) ( x + 3)

Directrix (directrix):

2y + 8 = 13x + 39 2y - 13x = 39 -8 2y - 13x = 31

Directrix (directrix):

In standard form, -13x + 2y = 31 The Standard Form for a linear equation in two variables, x and y, is usually given as Ax + By = C

Directrix (directrix):

Question?